OpenStudy (anonymous):

factor 343x^3+125x^6

OpenStudy (angela210793):

x^3(5x+7)*(25x^2-35x+49)

OpenStudy (anonymous):

343x^3+125x^6 =x^3[(7)^3+(5x)^3] =x^3(7+5x)((7)^2-(7)(5x)+(5x)^2) =x^3(7+5x)(49-35x+25x^2) =x^3(5x+7)(25x^2-35x+49)